OpenStudy (anonymous):

Use a half-angle formula to find the exact value of the given expression. sin π/8

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

\[\frac{\pi}{12}\] is half of \[\frac{\pi}{6}\] use that one

OpenStudy (anonymous):

so pi/6 + pi/6?